a number consist of two digits whose sum is 9.If 27 is subtracted form the number. Its digits are reversed
Answer by robertb(4012)

let the tens digit be x, y the ones digit. Then x+ y = 9, and
10x + y - 27 = 10y + x.
The second equation is equivalent to 9x - 9y = 27, or x - y = 3.
Combining this with the first equation, we get x = 6, and y = 3.
The original number is 63.